OpenStudy (anonymous):

f '(x) =-9 (cos(x)+x cos(x)+sin(x)-x sin(x)) solve for f '(-pi/4)

OpenStudy (anonymous):

you can plug this into the caluculator?

OpenStudy (agreene):

This is just a matter of plugging in, remember your unit circle: \[\cos \frac{-\pi }4=\frac{\sqrt{2}}{2}\] \[\sin \frac{-\pi }4=\frac{\sqrt{2}}2\] just plug and for your cos/sin and go go combines/simplifies.